OpenStudy (anonymous):

The current I in an electrical conductor varies inversely as the resistance R of the conductor. The current is 5 amperes when the resistance is 910 ohms. What is the current when the resistance is 550 ohms?

OpenStudy (anonymous):

if x and y vary inversely, they their product is a constant i.e. x*y = constant in first case x1 = 5, y1 = 910 so x1*y1 = 5*910 = 4550 in second case x2 = ???, y2 = 550 now x2*y2 = 4550 x2*550 = 4550 x2 = 4550/550 x2 = 8.27